FILTERS :
RILIS:
Time I
S
O
G
P
S
Comments
08:33 SU
Preparing for target change. #611 ->#609
Target already cold. Thanks Collections team for doing that.

Now stopping HT, Electrode out, transformer stop, water stop, etc.

/ES/SM
09:05 SU
This time we had to turn the GPS HT key in the HT room to obtain the Vacuum OFF interlock OK. As it used to be.
Last target change on HRS we did not need to trun the HT key. Vacuum OFF did work then..

/ES
09:23 SU
Target Number: #611 UC
GPSProtons countuAmps.hWindow TimeNumber Of ShotsShort 1.18921821102406118E1852.926 7d 22h 51min 21sec 131116Long 7.462661034185035E193321.2501 533d 12h 30min 28sec 3332889Vistar 1.18921821102406118E1852.926 7d 22h 51min 21sec 131116
name: Screenshot.jpg
desc:Screenshot of logbook
09:25 SU
GPS10 vented. Interlock shutter, etc. OK.
Ready for target change.

/ES
09:49 SU
Proposed heating values (Sebastian).

Target 600A
Line 220A (weekend) (255A later for tuning)

OVEN1 connected to the Sm
OVEN2 connected to the Nd (Neodymium)

Tunning stable using 39K and 142Nd (stable, OVEN2 massmarker) (run on 140Nd radioactive)

/ES
09:51 SU
Target #609 cabling connection done
name: 20170922095213.png
desc:
name: 20170922095733.png
desc:
10:03 SU
Target #611 unclamped now.

/ES
10:08 SU
Set the Beam Settings back to nominal (2uA max p-current, 3.5E13ppp max, Dipoles p-beam set to 0A).

/ES
name: 20170922100934.png
desc:
10:09 SU
Switched GPS10 to pumping.

/ES
10:29 SU
Target #611 stored on GPS shelf 411
Target #609 installed on the GPS Front End with the Montrac.

/GL
11:01 SU
Sector 6 / MEDICIS patrol done.
HT room patrol done.

Asked PSB to set the security chain.

Ready for protons in a few minutes (if wanted).

/ES/EF/SM
11:02 SU
GSP10 setcor pumped.
Cooling water on.
Transfomer & target/line power supplies on.

/ES
11:10 SU
Started target/line heating.

HT will need rearming in the HT room. GPS10 was not pumped yet when we did the patrol. Setting up will not start until Monday. Heating now.

/ES/SM
name: 20170922111155.png
desc:
14:18 SU
Started HT! Now at 30.2kV for Trap injection next week.

/ES
16:03 SU
Target at 600A, Line at 220A.

Set the electrode to 60mm.

/ES
16:45 SU
Increased the line from 220A to 255A.

We see a bit of 39K coming:

Total beam GPS.FC0600 11nA
GPS.FC490 9.3pA
GPS.FC558 5.2pA


See it on GPS.SC483 but not on GPS.SC556 (too little beam).
MAG70 calibrated.
HT 30.2kV
Electrode at 60mm
Settings from previous run at GPS

Did not see any 23Na.

/ES
name: 20170922164928.png
desc:
name: 20170922165118.png
desc:
name: 20170922170304.png
desc:
17:04 SU
Cooling down now the line back from 255A to 220A.

/ES